OpenStudy (anonymous):

Find the deriviative of y=x^2 - 3x - 1 at x=0

OpenStudy (agent47):

y'=2x-3 plug in x = 0

OpenStudy (anonymous):

y'=2x-3; y'(0)=2(0)-3=-3

OpenStudy (anonymous):

Im not entirely sure how you did that. Im supposed to find it using the formula of (f(x+h) - x)/h

OpenStudy (anonymous):

Oh. You haven't learned the shortcut yet. Using that, you have to expand it. For example, look at this: http://www.analyzemath.com/calculus/Differentiation/derivative_definition.html

OpenStudy (anonymous):

(H+x)^2 -3(x+H) - 1 -(X^2 -3X - 1) X^2 + 2hx + H^2 - 3x - 3h - 1 - x^2 + 3x + 1 2hx + H^2 -3h / h \[\prime= 2x\] Its supposed to be -3, what am I doing wrong?

OpenStudy (anonymous):

I think our teacher wants use to learn how to do it using this method before using any shortcuts.

OpenStudy (anonymous):

y=x^2 - 3x - 1 f(x+h)= (x+h)^2 - 3(x+h) - 1 f(x)= x^2 - 3x - 1 [((x+h)^2 - 3(x+h) - 1) - (x^2 - 3x - 1)]/ h

OpenStudy (anonymous):

[(x^2+2xh+h^2) - 3x+3h - 1 - x^2 + 3x +1]/h

OpenStudy (anonymous):

I see what I did wrong, 2hx + H^2 -3h / h turns into -3 Sorry for wasting you time, but I did learn a shortcut. Thanks!

OpenStudy (anonymous):

No problem.